suppose you sell a total of 292 tickets for a football game. an adult ticket cost $3.00 and a student ticket cost $1.00. The sponsors collect $470 in ticket sales. Write and solve a system to find the # of each ticket. I got this so far:
3a + 1s =470
help

Found 2 solutions by stanbon, xinxin:
Answer by stanbon(75887)

suppose you sell a total of 292 tickets for a football game. an adult ticket cost $3.00 and a student ticket cost $1.00. The sponsors collect $470 in ticket sales. Write and solve a system to find the # of each ticket.
---------------------
Quantity Equation: a + s = 292 tickets
Value Equation::: 1*a+3s = 470
----------
Substract and solve for "s":
2s = 178
s = 89 (# of student tickets sold)
------
Solve for "a":
a + s = 292
a + 89 = 292
s = 203 (# of adult tickets sold)
=================
Cheers,
Stan H.
=================

Answer by xinxin(76)

3a + 1s = 470 (1)
a + s = 292 (2)
You can arrange either (1) or (2) to get an equation with only one variable. For example, (2) can be rearranged as a = 292 - s (3). Next we need to combine (1) and (3) to get s: 3 * (292 - s) + s = 470, s = 203. So a is 89.